6 November 2007

RHIC Startup:
RHIC Shifts start this evening.
Cooldown complete. 
Temperature sensor and He-fill level sensor in 4OClock fixed.
Hi-pot problems cleared when valve box cleaned.  Testing at injection today.  Ramping up to the B-rho for deuterons plus 10%.  Doing a hysteresis ramp tonight.  Need to fix fans on Thursday.
X/Y arc swept and ready for beam.
Access Controls plan to finish today.
Start sweeping the ring after lunch today, starting at 12OClock, then 4, 6 and 8.
A5 pulse width too wide for 8 bunches of deuteron transfers.  Looking for RF solution with 6 bunch transfers.  KSS expects 4 days to make this work, and gives a 60% probability of success.  In this mode, filling deuterons would take 1 minute longer than normal, but this achieves the shortest bunches, best emittance, and highest intensity.
Inject Blue deuterons tonghit, with 1 of the four bunches at 50% intensity.  All other bunches at ~1.2E11 at booster input, 0.9E11 at AGS late.  Plan to begin test with kicker off to look for obstructions.  Then turn kicker on and inject again.  Permit system must be changed to allow injection into Bleu while Yellow still Off.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020